what is 2d+24+3d=84
Found 2 solutions by Alan3354, oberobic:
Answer by Alan3354(69443)

what is 2d+24+3d=84
--------
It's an equation.

Answer by oberobic(2304)

2d +24 +3d = 84
5d +24 = 84
5d = 60
d = 12